OpenStudy (anonymous):

Mark bought a plane ticket for $234, which was 80% of the original price. How much was the original price? Be sure to write your answer in dollars and cents.

OpenStudy (anonymous):

(234/80)x100=292.5 total amount is $292.5 20% of the amount=292.5-234=58.5

OpenStudy (texaschic101):

80% of x = 234 .80x = 234 x = 234/.80 x = 292.50 check... (.80)(292.50) = 234 234 = 234 (correct) ANSWER : original price is $ 292.50